# WebAgent-R1: Training Web Agents via End-to-End Multi-Turn Reinforcement Learning

### v1

@@ Appendix A Web Environment
WebArena (Zhou et al., 2024a) is a realistic, self-hostable web environment for developing LLM-based agents. It comprises 812 real-world web tasks spanning diverse domains, including social forum (Reddit), collaborative coding (GitLab), e-commerce content management system (CMS), open street map (Map), and online shopping (OneStopShop). WebArena-Lite (Liu et al., 2025) is a curated version of WebArena designed for more reliable evaluation. It selects 165 representative tasks for human verification as the evaluation set and uses the remaining 647 tasks for training. It also provides 9,460 trajectories automatically annotated by program-based solvers for behavior cloning. For each website, the authors (Liu et al., 2025) summarize the core functionalities and valid items and construct a set of task prototypes and manually implement rule-based solvers using Playwright scripts for each prototype. The corresponding solvers are executed on the websites to collect ground-truth trajectories. In total, this produces 1,186 valid training samples comprising 9,460 trajectories, released under the Apache License 2.0.
@@
<answer> do(action=‘‘Click’’, element=‘‘7’’) </answer>
@@ Appendix E Qualitative Examples
System Prompt You are a professional web browsing agent assistant that can fulfill user’s high-level instructions. Given simplified html of the browsed webpage at each step, you plan operations in python-style pseudo code using provided functions. You should first think about the reasoning process as an internal monologue and then decide an action. The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively, i.e., responding in the following format: <think> ... </think> <answer> ... </answer>. More details about the code action: Your action should be readable, simple. Please generate **ONLY ONE ACTION** in one round. Predefined functions are as follows:

### v2

@@ Appendix A Web Environment
WebArena (Zhou et al., 2024a) is a realistic, self-hostable web environment for developing LLM-based agents. It comprises 812 real-world web tasks spanning diverse domains, including social forum (Reddit), collaborative coding (GitLab), e-commerce content management system (CMS), open street map (Map), and online shopping (OneStopShop). WebArena-Lite (Liu et al., 2025a) is a curated version of WebArena designed for more reliable evaluation. It selects 165 representative tasks for human verification as the evaluation set and uses the remaining 647 tasks for training. It also provides 9,460 trajectories automatically annotated by program-based solvers for behavior cloning. For each website, the authors (Liu et al., 2025a) summarize the core functionalities and valid items and construct a set of task prototypes and manually implement rule-based solvers using Playwright scripts for each prototype. The corresponding solvers are executed on the websites to collect ground-truth trajectories. In total, this produces 1,186 valid training samples comprising 9,460 trajectories, released under the Apache License 2.0.
@@
<answer> do(action=‘‘Click", element=‘‘7") </answer>
@@ Appendix E Qualitative Examples
System Prompt You are a professional web browsing agent assistant that can fulfill user’s high-level instructions. Given simplified html of the browsed webpage at each step, you plan operations in python-style pseudo code using provided functions. You should first think about the reasoning process as an internal monologue and then decide an action. The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively, i.e., responding in the following format: <think> ... </think> <answer> ... </answer>.
More details about the code action: Your action should be readable, simple. Please generate **ONLY ONE ACTION** in one round. Predefined functions are as follows: